OpenStudy (briellshaw):

micheal is 7 his brother is 11 his oldest brother is 27 whhat is the diffrence between all of them?

OpenStudy (briellshaw):

i know the answer its 9

OpenStudy (mathstudent55):

What does that mean? I know what the difference between two numbers means, but I don't know what the difference between three numbers means.

OpenStudy (briellshaw):

27-11=16-7=9

OpenStudy (mikeymaximum):

Yes, I think thats right...